# The swap transpose on couplings translates to Petz’ recovery map on quantum channels

### v1

@@ 1.1. Motivation and main result
This note is concerned with a concept where quantum channels realize the transport, and quantum couplings are in a one-to-one correspondence with channels. This concept was introduced by De Palma and Trevisan [26, 27], and it is a quantum counterpart of the classical transport theory with quadratic cost — see also [11, 13, 12, 68] for further recent developments in this direction. We prove in an elementary way that, in this setting, two natural operations, the swap transpose on couplings, and taking the Petz recovery map [56, 55, 1] on channels, are two sides of the same coin. We phrase the precise statement in Theorem 1— all the necessary notions and notations will be introduced in Subsection 1.2. We note that Theorem 1 is similar in nature to a result of Duvenhage and Snyman [33] that relates taking the dual of a completely positive unital map to a certain involution on transport plans — see Propositions 3.4 and 3.9 and the conclusion eq. (11) there.
@@ 1.2. Basic notions, notation
Let $\rho\in\mathcal{S}\left(\mathcal{H}\right)$ be a state, and let $\Phi:\mathcal{T}_{1}(\mathrm{supp}(\rho))\to\mathcal{T}_{1}(\mathcal{H})$ be a quantum channel, that is, a completely positive and trace-preserving map. Then the Petz recovery map (see, e.g., [65, Section 12.3]) for $\rho$ and $\Phi$ is the completely positive and trace-preserving map $\Phi_{\text{rec}}^{\rho}$ defined by
@@
Assume that $\rho$ and $\omega$ admit the following spectral resolutions: $\rho=\sum_{i}\lambda_{i}|e_{i}\rangle\langle e_{i}|$ and $\omega=\sum_{j}\mu_{j}|g_{j}\rangle\langle g_{j}|.$ Let $\{f_{i}\}$ and $\{h_{j}\}$ denote the dual bases of $\{e_{i}\}$ and $\{g_{j}\},$ respectively, in $\mathcal{H}^{*}.$ Then
@@
$=\left(\Phi\otimes\mathrm{id}_{\mathcal{T}_{1}(\mathcal{H}^{*})}\right)\left(\sum_{i,j}\sqrt{\lambda_{i}\lambda_{j}}|e_{i}\rangle\langle e_{j}|\otimes\left(|e_{j}\rangle\langle e_{i}|\right)^{T}\right)=\sum_{i,j}\sqrt{\lambda_{i}\lambda_{j}}\Phi\left(|e_{i}\rangle\langle e_{j}|\right)\otimes\left(|e_{j}\rangle\langle e_{i}|\right)^{T}.$
@@
$|g_{k}\rangle\langle g_{l}|=\sum_{r,s}\langle e_{r}||g_{k}\rangle\langle g_{l}||e_{s}\rangle|e_{r}\rangle\langle e_{s}|\text{ and }|g_{l}\rangle\langle g_{k}|=\sum_{m,n}\langle e_{m}||g_{l}\rangle\langle g_{k}||e_{n}\rangle|e_{m}\rangle\langle e_{n}|.$
Accordingly,
@@
Acknowledgement.
We thank Melchior Wirth for his comments on a preliminary version of this manuscript, and in particular for introducing us to related results in the von Neumann algebraic setting.

### v2

@@ 1.1. Motivation and main result
This note is concerned with a concept where quantum channels realize the transport, and quantum couplings are in a one-to-one correspondence with channels. This concept was introduced by De Palma and Trevisan [26, 27], and it is a quantum counterpart of the classical transport theory with quadratic cost — see also [11, 13, 12, 69] for further recent developments in this direction. In this setting, we prove in an elementary way that two natural operations, the swap transpose on couplings, and taking the Petz recovery map [57, 56, 1] on channels, are two sides of the same coin. We phrase the precise statement in Theorem 1— all the necessary notions and notations will be introduced in Subsection 1.2.
It is important to note that finite-dimensional versions of our results, with different conventions and terminology, have already been established. In [32, Theorem 5.2], Duvenhage, Oerder, and van den Heuvel prove that the Accardi-Cecchini dual [1] of a completely positive map (on operators acting on a finite dimensional Hilbert space) coincides with the dual corresponding to swapping (without the transpose, in contrast to the convention we use). Furthermore, in [55], Parzygnat and Fullwood proved essentially the same finite-dimensional result in the context of Bayes’ rules in the quantum setting — see in particular eq. (14) and eq. (25) there.
Acknowledgement.
We thank Rocco Duvenhage and Arthur Parzygnat for drawing our attention to their closely related works [32] and [55] after the appearence of the first version of our manuscript on arXiv, and for their detailed explanations of the connections. We are also gratefut to Melchior Wirth for his comments on a preliminary version of this manuscript, and in particular for introducing us to related results in the von Neumann algebraic setting.
@@ 1.2. Basic notions, notation
Let $\rho\in\mathcal{S}\left(\mathcal{H}\right)$ be a state, and let $\Phi:\mathcal{T}_{1}(\mathrm{supp}(\rho))\to\mathcal{T}_{1}(\mathcal{H})$ be a quantum channel, that is, a completely positive and trace-preserving map. Then the Petz recovery map (see, e.g., [66, Section 12.3]) for $\rho$ and $\Phi$ is the completely positive and trace-preserving map $\Phi_{\text{rec}}^{\rho}$ defined by
@@
Assume that $\rho$ and $\omega$ admit the following spectral resolutions: $\rho=\sum_{i=1}^{\infty}\lambda_{i}|e_{i}\rangle\langle e_{i}|$ and $\omega=\sum_{j=1}^{\infty}\mu_{j}|g_{j}\rangle\langle g_{j}|.$ Let $\{f_{i}\}$ and $\{h_{j}\}$ denote the dual bases of $\{e_{i}\}$ and $\{g_{j}\},$ respectively, in $\mathcal{H}^{*}.$ Then
@@
$=\left(\Phi\otimes\mathrm{id}_{\mathcal{T}_{1}(\mathcal{H}^{*})}\right)\left(\sum_{i,j}\sqrt{\lambda_{i}\lambda_{j}}|e_{i}\rangle\langle e_{j}|\otimes\left(|e_{j}\rangle\langle e_{i}|\right)^{T}\right)=\sum_{i,j}\sqrt{\lambda_{i}\lambda_{j}}\Phi\left(|e_{i}\rangle\langle e_{j}|\right)\otimes\left(|e_{j}\rangle\langle e_{i}|\right)^{T},$
where by the infinite sum $\sum_{i,j}\sqrt{\lambda_{i}\lambda_{j}}|e_{i}\rangle\langle e_{j}|\otimes\left(|e_{j}\rangle\langle e_{i}|\right)^{T}$ we mean the trace norm limit
$\lim_{M,N\to\infty}^{\left|\left|\cdot\right|\right|_{1}}\sum_{i=1}^{M}\sum_{j=1}^{N}\sqrt{\lambda_{i}\lambda_{j}}|e_{i}\rangle\langle e_{j}|\otimes\left(|e_{j}\rangle\langle e_{i}|\right)^{T},$
and similarly,
$\sum_{i,j}\sqrt{\lambda_{i}\lambda_{j}}\Phi\left(|e_{i}\rangle\langle e_{j}|\right)\otimes\left(|e_{j}\rangle\langle e_{i}|\right)^{T}=\lim_{M,N\to\infty}^{\left|\left|\cdot\right|\right|_{1}}\sum_{i=1}^{M}\sum_{j=1}^{N}\sqrt{\lambda_{i}\lambda_{j}}\Phi\left(|e_{i}\rangle\langle e_{j}|\right)\otimes\left(|e_{j}\rangle\langle e_{i}|\right)^{T}.$
@@
$|g_{k}\rangle\langle g_{l}|=\sum_{r,s}\langle e_{r}||g_{k}\rangle\langle g_{l}||e_{s}\rangle|e_{r}\rangle\langle e_{s}|\text{ and }|g_{l}\rangle\langle g_{k}|=\sum_{m,n}\langle e_{m}||g_{l}\rangle\langle g_{k}||e_{n}\rangle|e_{m}\rangle\langle e_{n}|,$
where, again, the infinite sums stand for the trace-norm limit of the finite partial sums. Accordingly,